OpenStudy (anonymous):

find the exact value of sin(cos^-1(3/7)). a. undefined b. (2sqrt10)/7 c. 40/49 d. 7/(2sqrt10)

OpenStudy (xapproachesinfinity):

hmm let call \[\arccos(\frac{3}{7})=\theta\] so \[\cos\theta=\frac{3}{7}\] we can draw a picture that can help us here

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

we need to find ? we can use pythagorean theorem \[x^2+3^2=7^2 \Longrightarrow x=\sqrt{49-9}=2\sqrt{10}\]

OpenStudy (xapproachesinfinity):

x is the missing side we found it

OpenStudy (xapproachesinfinity):

now we do \[\sin \theta\] which is the same as \[\sin (\arccos(\frac{3}{7}))\]

OpenStudy (xapproachesinfinity):

i let you do this part...

OpenStudy (xapproachesinfinity):

from the triangle what is sin theta

OpenStudy (anonymous):

I got .903

OpenStudy (anonymous):

remember sin is opp/hyp

OpenStudy (anonymous):

which would be b correct

OpenStudy (xapproachesinfinity):

sin=opp/hyp like @magepker728 the opposite is \[2\sqrt{10}\] the hyp=7

OpenStudy (xapproachesinfinity):

i guess you are not interested lol